OpenStudy (anonymous):

The volume of a cylinder increased by 8 times the original volume. How many times greater are each of the dimensions of that new cylinder? A: 2 B: 3 C: 4 D: 8

OpenStudy (anonymous):

nvm I got the awnser done.

OpenStudy (saifoo.khan):

There's a ratio formula:\[\Large A = L^2\]And \[\Large V = L^3\]

OpenStudy (saifoo.khan):

Where A = area and v = volume. L are the sides.

OpenStudy (raden):

2, just guess...

OpenStudy (saifoo.khan):

Now you should use the second formula: